OpenStudy (anonymous):

Plans for an arena map for a popular video game are drawn on a coordinate grid. A new corridor of the arena lies on the line Y= 5/2x. A perpendicular corridor passes through the point (10,0). Write the equation of the line that contains the new corridor. A) y=5/2x-25 B) Y=-2/5x+4 C) Y=2/5x-4 D) Y=-5/2x+25

OpenStudy (anonymous):

doing the same test as you right now

OpenStudy (anonymous):

haha are you really?

OpenStudy (anonymous):

yeah I am

OpenStudy (charlotte123):

looks like its C :D

OpenStudy (anonymous):

That's what i'm getting out of it!

OpenStudy (charlotte123):

HUZZAH ┌(-_-) ┐└(-_-)┘┌(-_-) ┐└(-_-)┘

OpenStudy (charlotte123):

(10, 0) makes y 0 right? :O

OpenStudy (charlotte123):

oh wait my bad XD never mind :3 >.>

OpenStudy (anonymous):

B) Y=-2/5x+4 will be the answer

OpenStudy (anonymous):

I got b as well

OpenStudy (charlotte123):

OH WAIT ITS B Y=-2/5x+4 not - >.>

OpenStudy (charlotte123):

http://openstudy.com/study#/updates/526d9517e4b0ff17247677e6 derp .-.

OpenStudy (anonymous):

yes u gt it rt

OpenStudy (anonymous):

Thank you guys so much

OpenStudy (charlotte123):

:D Savior, was our pleasure :D as long as u get a rest from all daht peachin work ^_^

OpenStudy (anonymous):

hope you do good on the test

OpenStudy (anonymous):

aallll the bst

OpenStudy (charlotte123):

Good luck Savior 。(⌒∇⌒。)

OpenStudy (anonymous):

Which of the following equations is parallel to the line passing through A(-2, 7) B(-2,-2)? A) y=-2 B) Y= x C) 3x=5 D) -2x=7y